# GPSMamba: A Global Phase and Spectral Prompt-guided Mamba for Infrared Image Super-Resolution

### v1

@@ Introduction
×2$\times$ 2× 2 EDSR[CVPRW 2017] (Lim et al. 2017) 1,369 39.0493 11.8196 0.9414 39.8902 8.9865 0.9528 44.1770 2.7845 0.9713
@@
×4$\times$ 4× 4 EDSR[CVPRW 2017] (Lim et al. 2017) 1,369 34.5219 30.1273 0.8548 35.1740 23.9917 0.8723 40.1190 6.8819 0.9482
@@
[…]
@@
where $\odot$ denotes element-wise multiplication and ||$\cdot$|| ⋅ | computes the magnitude. This semantically-guided spectral matching compels our model to prioritize the fidelity of frequency components originating from regions identified as thermally salient.
@@
Datasets. To ensure a rigorous and fair comparison with SOTA methods, we adopt the experimental protocol established by recent works like IRSRMamba. Our model was trained on the widely-used M3FD dataset(Liu et al. 2022b). For evaluation, we benchmark our GPSMamba on three standard infrared test sets: result-A(Liu 2018; Huang et al. 2021), result-C(Zhang et al. 2017; Huang et al. 2021), and CVC10(Campo et al. 2012). Following standard practice, the LR inputs were generated by applying bicubic downsampling to the HR images for scale factors of ×2$\times$ 2× 2 and ×4$\times$ 4× 4.
@@
(a) Error Distribution Across Models on result-A (Scale ×4$\times$ 4× 4)
(b) Error Distribution Across Models on CVC10 (Scale ×4$\times$ 4× 4)
Figure 3: Global error distribution analysis for various super-resolution models on the result-A and CVC10 datasets at scaling factor of ×4$\times$ 4× 4. Each subfigure presents a comparative error breakdown across different ranges, quantifying the proportion of minimal, moderate, high, and severe reconstruction errors.
@@
Figure 4: Perceptual quality assessment of super-resolution models on result-A. Larger area indicates better perceptual performance. (Scale ×2$\times$ 2× 2)
We conduct a comprehensive quantitative comparison against a suite of SOTA methods, including recent top-conference models like CATANet, MambaOut, and the current leading architecture, IRSRMamba. The full results for ×2$\times$ 2× 2 and ×4$\times$ 4× 4 super-resolution on three benchmark datasets are presented in Table 1. Our proposed model, GPSMamba, demonstrates exceptional performance. For the standard ×2$\times$ 2× 2 SR task, it achieves either the best or second-best results across all datasets and metrics, establishing its competitive standing. For instance, on the result-C dataset, our method surpasses IRSRMamba in PSNR and matches its leading SSIM score. The superiority of our approach becomes more pronounced in the more challenging ×4$\times$ 4× 4 SR task, where recovering high-frequency details is significantly more difficult. As shown in Table 1, our GPSMamba consistently sets a new SOTA, outperforming all competing methods on all metrics across all three datasets. On the result-C dataset, we achieve a PSNR of 35.4007 dB, surpassing the previous best (IRSRMamba) by a significant margin of nearly 0.1 dB. This marked improvement underscores our model’s enhanced capability to reconstruct faithful and structurally coherent images from highly degraded inputs. To dissect performance beyond aggregate metrics, we analyze the pixel-wise error distribution in Fig. 3. The visualization for the ×4$\times$ 4× 4 task reveals our model’s advantage: it reconstructs 63% of pixels with minimal error ([0, 5)), surpassing all competitors (result-A (Scale ×4$\times$ 4× 4)). Consequently, it produces the fewest pixels in the high-error categories ([10, $\infty$)). To complement our fidelity-based analysis, we assess perceptual quality using four non-reference metrics. As illustrated in Fig. 4, our method’s superiority is most evident on the challenging CLIP-IQA metric, where it achieves a leading score of 0.3043 and surpasses the current SOTA, IRSRMamba (0.2887). This advantage, which points to a more semantically coherent and realistic reconstruction, is supported by highly competitive performance across DBCNN, BRISQUE, and NIQE. The radar chart visually synthesizes these findings, showing our model’s performance envelope (blue) achieves the most favorable overall balance.
Figure 5: Visual comparison for ×4$\times$ 4× 4 super-resolution on images fused2 (top) and result12 (bottom). Ours GPSMamba reconstructs significantly sharper details, a claim substantiated by the minimal reconstruction errors (darker blue) shown in the corresponding residual maps. Best view in zoom.
@@
Qualitative Results. As in Fig.5, qualitative comparison for ×4$\times$ 4× 4 super-resolution on the result-A/result-C. Our model, GPSMamba, excels in restoring sharp textures and object contours where other methods fail. In the fused2 scene (top), it renders crisp, legible text on the sign, a detail completely lost to blurring in competing results. Similarly, in result12 (bottom), it preserves the pedestrian’s sharp silhouette. This superior fidelity is further substantiated by the residual maps (right), where our method exhibits minimal reconstruction error (darker blue) in these critical high-frequency regions, confirming a more accurate reconstruction.

### v3

@@ Introduction
$\times 2$ EDSR[CVPRW 2017] (Lim et al. 2017) 1,369 39.0493 11.8196 0.9414 39.8902 8.9865 0.9528 44.1770 2.7845 0.9713
@@
$\times 4$ EDSR[CVPRW 2017] (Lim et al. 2017) 1,369 34.5219 30.1273 0.8548 35.1740 23.9917 0.8723 40.1190 6.8819 0.9482
@@
where $F_{FFT}(\cdot)$ represents the 2D FFT and $\text{angle}(\cdot)$ extracts the phase component. Since image phase is known to carry critical information about structural layout, this loss term is vital for preserving sharp edges, object boundaries, and overall structural integrity in the reconstructed output.
@@
where $\odot$ denotes element-wise multiplication and $|\cdot|$ computes the magnitude. This semantically-guided spectral matching compels our model to prioritize the fidelity of frequency components originating from regions identified as thermally salient.
@@
Datasets. To ensure a rigorous and fair comparison with SOTA methods, we adopt the experimental protocol established by recent works like IRSRMamba. Our model was trained on the widely-used M3FD dataset(Liu et al. 2022b). For evaluation, we benchmark our GPSMamba on three standard infrared test sets: result-A(Liu 2018; Huang et al. 2021), result-C(Zhang et al. 2017; Huang et al. 2021), and CVC10(Campo et al. 2012). Following standard practice, the LR inputs were generated by applying bicubic downsampling to the HR images for scale factors of $\times 2$ and $\times 4$.
@@
(a) Error Distribution Across Models on result-A (Scale $\times 4$)
(b) Error Distribution Across Models on CVC10 (Scale $\times 4$)
Figure 3: Global error distribution analysis for various super-resolution models on the result-A and CVC10 datasets at scaling factor of $\times 4$. Each subfigure presents a comparative error breakdown across different ranges, quantifying the proportion of minimal, moderate, high, and severe reconstruction errors.
@@
Figure 4: Perceptual quality assessment of super-resolution models on result-A. Larger area indicates better perceptual performance. (Scale $\times 2$)
We conduct a comprehensive quantitative comparison against a suite of SOTA methods, including recent top-conference models like CATANet, MambaOut, and the current leading architecture, IRSRMamba. The full results for $\times 2$ and $\times 4$ super-resolution on three benchmark datasets are presented in Table 1. Our proposed model, GPSMamba, demonstrates exceptional performance. For the standard $\times 2$ SR task, it achieves either the best or second-best results across all datasets and metrics, establishing its competitive standing. For instance, on the result-C dataset, our method surpasses IRSRMamba in PSNR and matches its leading SSIM score. The superiority of our approach becomes more pronounced in the more challenging $\times 4$ SR task, where recovering high-frequency details is significantly more difficult. As shown in Table 1, our GPSMamba consistently sets a new SOTA, outperforming all competing methods on all metrics across all three datasets. On the result-C dataset, we achieve a PSNR of 35.4007 dB, surpassing the previous best (IRSRMamba) by a significant margin of nearly 0.1 dB. This marked improvement underscores our model’s enhanced capability to reconstruct faithful and structurally coherent images from highly degraded inputs. To dissect performance beyond aggregate metrics, we analyze the pixel-wise error distribution in Fig. 3. The visualization for the $\times 4$ task reveals our model’s advantage: it reconstructs 63% of pixels with minimal error ([0, 5)), surpassing all competitors (result-A (Scale $\times 4$)). Consequently, it produces the fewest pixels in the high-error categories ([10, $\infty$)). To complement our fidelity-based analysis, we assess perceptual quality using four non-reference metrics. As illustrated in Fig. 4, our method’s superiority is most evident on the challenging CLIP-IQA metric, where it achieves a leading score of 0.3043 and surpasses the current SOTA, IRSRMamba (0.2887). This advantage, which points to a more semantically coherent and realistic reconstruction, is supported by highly competitive performance across DBCNN, BRISQUE, and NIQE. The radar chart visually synthesizes these findings, showing our model’s performance envelope (blue) achieves the most favorable overall balance.
Figure 5: Visual comparison for $\times 4$ super-resolution on images fused2 (top) and result12 (bottom). Ours GPSMamba reconstructs significantly sharper details, a claim substantiated by the minimal reconstruction errors (darker blue) shown in the corresponding residual maps. Best view in zoom.
@@
Qualitative Results. As in Fig.5, qualitative comparison for $\times 4$ super-resolution on the result-A/result-C. Our model, GPSMamba, excels in restoring sharp textures and object contours where other methods fail. In the fused2 scene (top), it renders crisp, legible text on the sign, a detail completely lost to blurring in competing results. Similarly, in result12 (bottom), it preserves the pedestrian’s sharp silhouette. This superior fidelity is further substantiated by the residual maps (right), where our method exhibits minimal reconstruction error (darker blue) in these critical high-frequency regions, confirming a more accurate reconstruction.